OpenStudy (anonymous):

what is the simplified form of the expression (m^-1m^5/m^-2)^-3

OpenStudy (anonymous):



jigglypuff314 (jigglypuff314):

when multiplying same variables that have exponents, the exponents add so for example, x^-2 x^7 = x^(-2+7) = x^5

OpenStudy (anonymous):

so (m^4 /m^-2)^-3

jigglypuff314 (jigglypuff314):

yep :) then when dividing same variables that have exponents the exponents subtract so for example, x^5 / x^-3 = x^(5-(-3)) = x^(5+3) = x^8

OpenStudy (anonymous):

(m^6)^-3

jigglypuff314 (jigglypuff314):

yep :) then when a variable with an exponent is raised to another exponent, the exponents multiply so for example, (x^2)^-4 = x^(2*(-4)) = x^-8

OpenStudy (anonymous):

m^-18

jigglypuff314 (jigglypuff314):

yep :) then there is a general rule that teachers don't like negative exponents so when there are negative exponents, they flip... so like x^-1 = 1/ x^1 and 1/x^-2 = x^2

OpenStudy (anonymous):

Okay so the answer would be 1/m^18?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

Thank you so much that was soooo helpful!!! ::D

jigglypuff314 (jigglypuff314):

You're Welcome :)